solve05 Question (5 points) See page 100 One of the most common functional forms used in economics is Cobb-Douglas. A Cobb-Douglas
Answer #1

462, 42= xay l-x 1) 20 Ох ao = xich-lyld a mun δυ (1-x) xx y 1-X-7 = muy oy MRS = more muy xxx-lyl-x (1-x) xx y1d-1 = - x x g- (3) Prezet Pyy = m i Pekt Bo | 1-8) P2 x = m. a ZTP & PR 2 (1-x) Pre-2 = m PRK [&t 127 - xm. Р.ҡ = am xCP2, py, my = am P2(4) Share of income spent on a = Prein m = Beam PK m share of income spent ony = Pyy m. = Py (2-x) m m (1-2).m ma = 11-0) (22